### v2

# Limitations of the decoding-to-LPN reduction via code smoothing

Abstract.
The Learning Parity with Noise (LPN) problem underlies several classic cryptographic primitives. Researchers have endeavored to demonstrate the algorithmic difficulty of this problem by attempting to find a reduction from the decoding problem of linear codes, for which several hardness results exist. Earlier studies used code smoothing as a technical tool to achieve such reductions, showing that they are possible for codes with vanishing rate. This has left open the question of attaining a reduction with positive-rate codes. Addressing this case, we characterize the efficiency of the reduction in terms of the parameters of the decoding and LPN problems. As a conclusion, we isolate the parameter regimes for which a meaningful reduction is possible and the regimes for which its existence is unlikely.
The authors are with the Department of ECE and Institute for Systems Research, University of Maryland, College Park, MD 20817. Emails: {madhura,abarg}@umd.edu.
1. Introduction
The Learning Parity with Noise (LPN) problem underlies several classic cryptographic primitives, including symmetric encryption [12, 14], public-key cryptography [1], and collision-resistant hashing [6, 24]. For this reason, the problem of computational complexity of LPN has attracted much attention in the literature. One line of work to show the hardness of this problem suggests to construct a reduction from the decoding problem of linear codes, arguing therefore that LPN is at least as hard as decoding generic codes. Reductions of this kind known to date [6, 23, 8] rely on the technique known as code smoothing. To establish their claims, the authors of these works have to make certain assumptions about the parameters of code sequences involved in the constructions. Specifically, meaningful reductions have been established only for code sequences of asymptotically vanishing rate, whereas the regime of positive rates so far has proved beyond reach. While recent works have focused on improved reductions, in this work we address a different possibility, namely that such reductions are not possible.
The LPN problem was proposed in [3] in a cryptographic context and was the subject of many subsequent works. Algorithms for solving it were designed in [4, 16] among others. In the search (or computational) version of the LPN problem, the goal is to recover a uniform random secret $m\in{\mathbb{F}}_{2}^{k}$ with high probability from ${N}$ samples of the form $(a_{i},a_{i}^{\intercal}m+b_{i})_{i=1}^{N},$ where the vectors $a_{i}$ are sampled uniformly and independently from ${\mathbb{F}}_{2}^{k}$ and $b_{i}\sim\operatorname{Ber}(\delta)$ are independent Bernoulli random bits. We call the parameter $\delta$ the noise rate of the corresponding LPN problem.
Studies into the hardness of LPN are motivated by an analogous computational task in the domain of lattice cryptography called Learning with Errors (LWE) [21, 17]. In this problem, one wants to recover a secret $m$ from a set of samples of the form $(a_{i},a_{i}^{\intercal}m+b_{i})$, where the vectors $(a_{i})_{i}$ are sampled from a large enough finite field ${\mathbb{F}}_{q}^{n}$, while $(b_{i})_{i}$ are sampled from a discrete Gaussian distribution. The primary technique employed in reducing worst-case lattice problems to LWE is lattice smoothing [17, 7]. By analogy, it is reasonable to expect that the worst-case decoding problems could similarly be reduced to the average-case LPN problem based on code smoothing, thereby supporting the hardness assumption.
Code smoothing can be informally described as follows. Given a code ${\mathscr{C}}\in{\mathbb{F}}_{2}^{n}$, we define a probability distribution $P_{\mathscr{C}}$ on ${\mathbb{F}}_{2}^{n}$ by normalizing the indicator function ${\mathbbm{1}}_{\mathscr{C}}$. Applying a noise operator $T$ to $P_{{\mathscr{C}}}$, we obtain another distribution. We say that $T$ smooths the code ${\mathscr{C}}$ if $TP_{\mathscr{C}}$ is close to the uniform distribution on ${\mathbb{F}}_{2}^{n}$. A more general version of this problem seeks to approximate any given “output” probability distribution $P$ and is known in information theory as channel resolvability [11], with numerous applications in information-theoretic problems. We refer to [18] for an overview of the literature as well as some recent results on code smoothing outside the context of this paper.
For a given $[n,k,d]$ linear code ${\mathscr{C}}$ with generator matrix ${\mathsf{G}}\in{\mathbb{F}}_{2}^{k\times n}$ and an integer $w\leq n$, the worst-case decoding problem involves identifying the message $m$ from a noisy codeword ${\mathsf{G}}^{\intercal}m+e$, with the promise that the Hamming weight of the error satisfies $|e|=w$ (we use column vectors throughout). The hardness of this problem was discussed in [9], and it was further shown in [10] it is NP-hard under the condition $w>d/2$. The non-promise version of the decoding problem is NP-hard even if one seeks to approximate the exact solution [2].
Before proceeding, let us make our expectations of the decoding-to-LPN reduction more explicit. Clearly, if the noise level $\delta=1/2,$ the LPN problem is not solvable. At the same time, symmetric cryptography becomes possible if $\delta=1/2-1/\text{poly}(k)$ [6], so we call a reduction meaningful if it yields an LPN problem with this or smaller noise level.
The general idea behind the reduction of decoding to LPN using smoothing is as follows. For a given linear code ${\mathscr{C}}$ with generator matrix ${\mathsf{G}}$, and an unknown error vector $e$ of weight $w$, the goal is to find a distribution $P$ on ${\mathbb{F}}_{2}^{n}$ so that sampling vectors $Z$ out of $P$ yields nearly uniform vectors ${\mathsf{G}}Z$, while $e^{\intercal}Z$ is a biased Bernoulli random variable that is nearly independent of ${\mathsf{G}}Z$. As will be seen below in the main text, finding the distribution $P$ such that $P_{{\mathsf{G}}Z}\approx P_{U_{k}}$ is equivalent to the smoothing. Previous works employed certain explicit distributions for generating the random vectors $Z$ and used Fourier-analytic arguments or the leftover hash lemma [13] to demonstrate the effectiveness of these choices.
The first paper devoted to the decoding-to-LPN reduction problem was the work of Brakersky et al. [6], followed by Yu and Zhang [23]. Using code smoothing, these papers show that LPN is at least as hard as solving a worst-case decoding problem. To claim this result, their authors have to make certain assumptions (explicit or implicit) that lead to limitations on the parameters of codes in the decoding part of the reduction. In particular, the reductions in [6, 23] become efficient once we require that the rate $k/n$ of the codes vanish as $n$ increases. Additionally the results in [6] rely on the assumption that the Hamming weights of the non-zero codewords are located between $l$ and $n-l$ for some $l$ (the authors call such codes balanced). The results of [23] require either balanced codes or certain independence conditions for the coordinates of the code. More recently, Debris-Alazard and Resch [8] presented a more general framework for utilizing code smoothing to reduce the decoding problem to the LPN problem. Their proof relies upon the work in [7] on lattice and code smoothing and does not involve any restrictions on the weights of the code, but their reduction is also limited by the vanishing rate assumption.
It remained unclear whether the limitation of zero rate stemmed from the particular smoothing distribution used in the cited works or from inherent issues within the smoothing approach itself. For example, in all prior studies [6, 23, 8], the authors employed a single universal smoothing distribution for all codes with the same parameters. However, in the reduction, this limitation is not essential; instead, distributions can vary depending on the specific codes being considered. Addressing this problem, we prove that for constant-rate codes, achieving a meaningful reduction with constant relative error weight $w/n$ is impossible regardless of the smoothing distribution. This result relies on the following qualification: we require that the vectors ${\mathsf{G}}Z$ (syndromes of the dual code) closely approximate uniformity, more formally, that the distance between their distribution and the uniform distribution on the space declines faster than $\operatorname{poly}(k)$. Allowing a slower approach to uniformity makes the reduction possible, although the parameter regime associated with it degrades the performance of the decoder implied by the LPN solver; see Sec. 2.2 for a more detailed discussion.
Our results can be interpreted as follows. The potential reduction depends on three parameters, namely the success probability of the LPN solver $\alpha$, the sample complexity of the solver $N$, and the smoothing parameter $\varepsilon.$ The success probability of the decoding task accomplished via the smoothing reduction is then expressed as $\alpha-N\varepsilon$. If the sample complexity scales as $N=2^{n^{c}},0<c<1$, then $\varepsilon$ has to be small to support the decoding. In this regime, we show that meaningful reductions for constant-rate codes are impossible. On the other hand, if $N$ has slower growth, then a reduction is possible, but the decoder performance is degraded.
A line of research closely related to the LPN problem addresses average-case decoding of linear codes. Two variants of the average-case decoding problem have been studied in the literature [8, 5]. Given a rnadom linear code and a fixed error weight, the search version of the average-case decoding problem seeks to recover the original codeword from a given noisy codeword with high probability. The decision version of this problem seeks to resolve whether a given vector is drawn from a uniform distribution or from a distribution that results from adding noise to a uniform random codeword. The authors of [8] reduced the worst-case decoding problem to the search version of the average-case decoding problem. Similarly, in [5], a reduction was obtained from the worst-case decoding problem to the decision version of the average-case decoding problem. In both cases, a crucial step was to reduce the worst-case decoding problem to the LPN problem, with code smoothing playing a pivotal role. Therefore, both approaches achieved meaningful reductions only for zero-rate codes. Since both results depend on reductions from worst-case decoding to LPN via code smoothing, our work suggests that generalizing them to positive-rate codes faces conceptual challenges and is unlikely.
2. The worst-case decoding problem and the LPN problem
2.1. Notation
We consider linear codes ${\mathscr{C}}\subset{\mathbb{F}}_{2}^{n}$ of length $n$. For a random vector $Z$ on ${\mathbb{F}}_{2}^{n}$ we denote by $P_{Z}$ its probability mass function (pmf). Given a pmf $P$ on ${\mathbb{F}}_{2}^{n}$, we write $Z\sim P$ to indicate that $Z$ is distributed according to $P$; hence $P_{Z}=P$. In particular, $P_{U_{n}}$ denotes the uniform distribution on ${\mathbb{F}}_{2}^{n}$, where $U_{n}$ is the uniform random vector on ${\mathbb{F}}_{2}^{n}$. For a code ${\mathscr{C}}\subset{\mathbb{F}}_{2}^{n}$, $P_{{\mathscr{C}}}$ denotes the uniform distribution on it and $X_{{\mathscr{C}}}$ is a uniform random codeword of ${\mathscr{C}}$. The notation $P_{\operatorname{Ber}(\delta)}$ refers to the pmf of a Bernoulli random variable $\xi$ with $P_{\xi}(1)=\delta$. The bias of $\xi$ is defined as $1/2-\delta$.
Given two probability measures $P$ and $Q$ defined on ${\mathbb{F}}_{2}^{n}$, define the total variation distance (TV-distance) as
(1) $d_{\text{\rm TV}}(P,Q)=\max_{A\subset{\mathbb{F}}_{2}^{n}}|P(A)-Q(A)|.$
It is well known that $d_{\text{\rm TV}}(P,Q)$ can be written as
(2) $d_{\text{\rm TV}}(P,Q)=\frac{1}{2}\sum_{x\in{\mathbb{F}}_{2}^{n}}|P(x)-Q(x)|.$
Following the usage in previous works, we say that a function is negligible, $f(k)=\operatorname{negl}(k)$, if $f(k)=o(k^{-a})$ for all $a>0$ as $k\to\infty$.
Let us formally define the worst-case decoding problem and the LPN problem.
Definition 2.1.
The worst-case decoding problem $\operatorname{WDP}(n,k,w)$ is defined as follows: Given
  1. (1)
    a matrix ${\mathsf{G}}\in{\mathbb{F}}_{2}^{k\times n}$
  2. (2)
    a vector $y\in{\mathbb{F}}_{2}^{n}$ of the form $y={\mathsf{G}}^{\intercal}m^{\prime}+e^{\prime}$ for some $m^{\prime}\in{\mathbb{F}}_{2}^{k}$ and $e^{\prime}\in{\mathbb{F}}_{2}^{n}$ with $|e^{\prime}|=w$,
find $m$ such that $y=G^{\intercal}m+e$ for some $e\in{\mathbb{F}}_{2}^{n}$ with $|e|=w$.
Definition 2.2.
The LPN problem $\operatorname{LPN}(k,\delta,{N},\alpha)$ with noise rate $\delta\in(0,1/2)$, sample complexity ${N}$, and success probability $\alpha$ is defined as follows: Given a collection of samples $(a_{i},a_{i}^{\intercal}m+b_{i})_{i=1}^{N}$, $a_{i},m\in{\mathbb{F}}_{2}^{k},b_{i}\in{\mathbb{F}}_{2}$, where
  1. (1)
    $m\sim P_{U_{k}}$ is fixed across all samples
  2. (2)
    $(a_{i},b_{i})\sim P_{U_{k}}P_{\operatorname{Ber}(\delta)}$ are chosen independently for each sample,
find $\hat{m}$ with $\Pr(\hat{m}=m)\geq\alpha$.
2.2. WDP-to-LPN reduction
In this section, we discuss the reduction from the worst-case decoding problem to LPN both in terms of the procedure that implements it and the assumptions for the parameters of the problem needed to prove hardness of LPN based on decoding. The reduction procedure presented below is due to [6].
Let $\mathcal{A}$ be an algorithm that solves $\operatorname{LPN}(k,\delta,{N},\alpha)$ in time $T$. Our goal is to show that this algorithm solves the decoding problem. Let ${\mathsf{G}}$ be a $k\times n$ generator matrix ${\mathsf{G}}$ of a code ${\mathscr{C}}$ and a vector (noisy codeword) ${\mathsf{G}}^{\intercal}m+e$, where the noise vector $e$ satisfies $|e|=w$ In order to find $m$ we proceed as follows:
  1. (a)
    Select a random vector $m^{\prime}$ uniformly from ${\mathbb{F}}_{2}^{k}$.
  2. (b)
    Find a distribution $P$ on ${\mathbb{F}}_{2}^{n}$ and a parameter $\varepsilon>0$ such that111 The probability distribution $P_{{\mathsf{G}}Z,e^{\intercal}Z}$ is given by $P_{{\mathsf{G}}Z,e^{\intercal}Z}(x,y)=\sum_{\begin{subarray}{c}z:{\mathsf{G}}z=x,\\ e^{\intercal}z=y\end{subarray}}P_{Z}(z),$ and similarly for other distributions of this kind.
    (3) $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,e^{\intercal}Z},P_{U_{k}}P_{\operatorname{Ber}(\delta)})\leq\varepsilon,$
    where $Z\sim P$. .
  3. (c)
    Generate samples $\{Z_{i}\}_{i=1}^{N}$ independently from $P$, and define $a_{i}={\mathsf{G}}Z_{i}$ and $b_{i}=e^{\intercal}Z_{i}$. Note that
    $Z_{i}^{\intercal}({\mathsf{G}}^{\intercal}m^{\prime}+{\mathsf{G}}^{\intercal}m+e)=a_{i}^{\intercal}(m+m^{\prime})+b_{i}.$
  4. (d)
    The set of pairs $(a_{i},a_{i}^{\intercal}(m+m^{\prime})+b_{i})_{i=1}^{N}$ is submitted as input to Algorithm $\mathcal{A}$.
  5. (e)
    If ${N}\varepsilon<\alpha$, Algorithm $\mathcal{A}$ outputs $m+m^{\prime}$ with success probability at least $\alpha-{N}\varepsilon$ in time $T$.
  6. (f)
    In conclusion, with probability $\alpha-{N}\varepsilon$ the message $m$ is found in time $T\!\cdot\operatorname{poly}(n,k)$.
Note that WDP is guaranteed to have a unique solution only if $w\leq(\text{dist}({\mathscr{C}})-1)/2$. If it does not, then the LPN problem obtained from it as described above recovers some solution to WDP with probability $\alpha-N\varepsilon$. Note also that the parameters $\delta,N,\alpha$ of the LPN problem are so far not related to $w$ and $n$. The interdependence of these parameters will be introduced through properties (4a)-(4b) of the reduction discussed below.
Hypothetically, if we could find $P$ such that for a vector $Z$ with $P_{Z}=P$, $P_{{\mathsf{G}}Z,e^{\intercal}Z}=P_{U_{k}}P_{\operatorname{Ber}(\delta)}$, it would imply that ${\mathsf{G}}Z$ and $e^{\intercal}Z$ are distributed as a product of a uniform distribution over ${\mathbb{F}}_{2}^{k}$ and a Bernoulli distribution. This ideal scenario would allow us to generate perfect inputs $(a_{i},a_{i}(m+m^{\prime})+b_{i})_{i}$ for the LPN solver, ensuring that $(a_{i})_{i}$ are uniformly distributed and independent from $(b_{i})_{i}$. Since this requirement generally cannot be satisfied, we relax this condition to (3). This relaxation degrades the performance of the decoder from $\alpha$ to $\alpha-{N}\varepsilon$. This follows by combining (3) with the definition of the total variation distance (1). Indeed, let ${\mathcal{E}}$ be the success event of the LPN solver, then the difference $|P_{{\mathsf{G}}Z,e^{\intercal}Z}({\mathcal{E}})-P_{U_{k}}P_{\operatorname{Ber}(\delta)}({\mathcal{E}})|\leq\varepsilon$, and the conclusion in Step (e) follows by the union bound.
The usefulness of a particular instantiation of the reduction depends on two parameters, the distance $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,e^{\intercal}Z},P_{U_{k}}P_{e^{\intercal}Z})$ and the bias of the bit $e^{\intercal}Z$.
First, an efficient algorithm for LPN results in an efficient probabilistic decoding procedure for a generic linear code. If LPN can be solved with a low-complexity procedure and if $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,e^{\intercal}Z},P_{U_{k}}P_{e^{\intercal}Z})$ is small, then the decoder succeeds with essentially the same probability as the LPN solver.
At the same time, the smaller the bias, the harder the LPN problem becomes. Indeed, if $\operatorname{bias}(e^{T}Z)=0$, then the observation is pure noise, and so the LPN problem is essentially unsolvable. Judging by this boundary case, we can expect that if $\operatorname{bias}(e^{\intercal}Z)=\operatorname{negl}(k)$ the problem is computationally hard, and hence of less theoretical interest. Furthermore, there are no known cryptographic implications for this parameter regime, while $\operatorname{bias}(e^{\intercal}Z)=O(1/\operatorname{poly}(k))$ implies symmetric cryptography [6].
Therefore, a meaningful reduction should satisfy the following two conditions:
(4a) $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})<\alpha/{N},$
(4b) $\operatorname{bias}(e^{\intercal}Z)=\Omega\Big{(}\frac{1}{\operatorname{poly}(k)}\Big{)}.$
Here condition (4a) essentially requires that random syndromes of the dual code ${\mathscr{C}}^{\bot}$ be approximately uniform. It arises because
$d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})\leq d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,e^{\intercal}Z},P_{U_{k}}P_{e^{\intercal}Z})\leq\varepsilon<\alpha/{N},$
where the first inequality holds because for any $X,Y,Z$,
$d_{\text{\rm TV}}(P_{XY},P_{ZY})=\frac{1}{2}\sum_{x,y}|P_{XY}(x,y)-P_{ZY}(x,y)|\\ \geq\frac{1}{2}\sum_{x}\Big{|}\sum_{y}(P_{XY}(x,y)-P_{ZY}(x,y))\Big{|}=d_{\text{\rm TV}}(P_{X},P_{Z}),$
and the last inequality holds because the probability of a successful reduction $\alpha-{N}\varepsilon$ must be strictly positive.
The central question is whether objectives (4a)-(4b) can be achieved simultaneously. In this paper, we show that the answer to this question depends on how closely the syndromes approximate uniformity. In Section 3.1, we show that if $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})=\operatorname{negl}(k)$, then for codes of positive rate and for constant relative error weight, the answer is negative, namely $\operatorname{bias}(e^{\intercal}Z)$ declines faster than any polynomial in $k$. We further show that to satisfy condition (4b) with $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})=\operatorname{negl}(k)$ the error weight necessarily has to be small, namely $O(\log k)$. This shows that for codes of constant rate and small error weight, the decoding problem can be efficiently reduced to the LPN problem. Of course, since decoding from a low-weight error is easy, this reduction does not contribute to the hardness proof of LPN.
With constant error weight, the only scenario where we can expect a meaningful reduction is when $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})=\Omega(1/\operatorname{poly}(k))$. This case is addressed in Sec. 3.2 where we show that such a reduction is achievable. At the same time, this parameter regime has an adverse effect on the accuracy of the decoder.
2.3. Fourier transform and Krawtchouk polynomials
The Fourier transform of a function $f:{\mathbb{F}}_{2}^{n}\rightarrow{\mathbb{R}}$ is defined as follows:
(5) $\widehat{f}(y)=\frac{1}{2^{n}}\sum_{x\in{{\mathbb{F}}_{2}^{n}}}f(x)(-1)^{x^{\intercal}y},\quad y\in{\mathbb{F}}_{2}^{n}.$
The inversion formula is given by
(6) $f(x)=\sum_{y\in{{\mathbb{F}}_{2}^{n}}}\widehat{f}(y)(-1)^{x^{\intercal}y},\quad x\in{\mathbb{F}}_{2}^{n}.$
We will also use the Fourier transform of the convolution $(f\ast g)(x):=\sum_{y}f(y)g(x-y)$, given by
(7) $(\widehat{f\ast g})({y})=2^{n}\widehat{f}({y})\widehat{g}({y})$
and Parseval’s formula
(8) $\frac{1}{2^{n}}\sum_{x}f(x)g(x)=\sum_{y}\widehat{f}({y})\widehat{g}({y}).$
Denote by $S(x,w)$ the sphere centered at $x$ with radius $w$ in the Hamming space. The Fourier transform of the indicator function of the sphere is given by $\widehat{\mathbbm{1}}_{S(0,w)}=\frac{1}{2^{n}}K_{w},$ where $K_{w}(x)={K}_{w}^{(n)}(x)=\sum_{j=0}^{w}(-1)^{j}\binom{|x|}{j}\binom{n-|x|}{w-j}$ is a Krawtchouk polynomial of degree $w$. Conversely, we have
(9) $\widehat{K}_{w}(x)={\mathbbm{1}}_{S(0,w)}(x),\quad x\in{\mathbb{F}}_{2}^{n}.$
If a function $f:{\mathbb{F}}_{2}^{n}\rightarrow{\mathbb{R}}$ depends only on the Hamming weight of its argument, we call it radial. Instead of ${\mathbb{F}}_{2}^{n},$ we sometimes write the domain of a radial function as $\{0,1,\dots,n\}$. With some abuse of notation, we also write $f(|x|)$ instead of $f(x)$, for instance, $K_{w}(i)$ rather than $K_{w}(x)$ if $|x|=i$.
Several bounds on the magnitude of $K_{t}(i)$ are known in the literature. We will use the following simple universal estimate.
Lemma 2.1.
[20, Lemma 2.4] There exist constants $c\in(0,1)$ and $C\geq 1$ such that for all $0\leq w\leq cn$ and $20\leq i\leq n/$ 2,
(10) $\frac{|K_{w}(i)|}{\binom{n}{w}}\leq C\Big{(}1-\frac{2w}{n}\Big{)}^{i}.$
For $n\geq 300$ it suffices to take $C=1$ and $c=0.16$.
2.4. Code smoothing
Here, we introduce the primary technique used in the reduction from decoding to LPN, known as code smoothing.
Definition 2.3.
We say that a random vector $Z$ (or distribution $P_{Z}$) $\varepsilon$-smooths a code ${\mathscr{C}}$ if $d_{\text{\rm TV}}(P_{X_{\mathscr{C}}+Z},P_{U_{n}})\leq\varepsilon$.
A convenient way of characterizing the $\varepsilon$-smoothing of an $[n,k]$ linear code ${\mathscr{C}}$ is given in the next lemma.
Lemma 2.2.
A random vector $Z$ (or a distribution $P_{Z}$) $\varepsilon$-smooths a code ${\mathscr{C}}$ if and only if
$d_{\text{\rm TV}}(P_{{\mathsf{H}}Z},P_{U_{n-k}})\leq\varepsilon,$
where ${\mathsf{H}}$ is the parity check matrix of ${\mathscr{C}}$.
Proof.
By a straightforward calculation,
$P_{{\mathsf{H}}Z}(u)=P_{{\mathsf{H}}(X_{\mathscr{C}}+Z)}(u)=\sum_{y:{\mathsf{H}}y=u}P_{X_{\mathscr{C}}+Z}(y)=2^{k}P_{X_{\mathscr{C}}+Z}(y_{u}),$
where $y_{u}\in{\mathbb{F}}_{2}^{n}$ is any vector satisfying ${\mathsf{H}}y_{u}=u$.
To prove the lemma we will show that
(11) $d_{\text{\rm TV}}(P_{{\mathsf{H}}Z},P_{U_{n-k}})=d_{\text{\rm TV}}(P_{X_{\mathscr{C}}+Z},P_{U_{n}}).$
Indeed,
$2d_{\text{\rm TV}}(P_{{\mathsf{H}}Z},P_{U_{n-k}})$ $=\sum_{u\in{\mathbb{F}}_{2}^{n-k}}\Big{|}P_{{\mathsf{H}}(X_{\mathscr{C}}+Z)}(u)-\frac{1}{2^{n-k}}\Big{|}$
$=\sum_{u\in{\mathbb{F}}_{2}^{n-k}}\Big{|}2^{k}P_{X_{\mathscr{C}}+Z}(y_{u})-\frac{1}{2^{n-k}}\Big{|}$
$=\frac{1}{|{\mathscr{C}}|}\sum_{y\in{\mathbb{F}}_{2}^{n}}\Big{|}2^{k}P_{X_{\mathscr{C}}+Z}(y)-\frac{1}{2^{n-k}}\Big{|}$
$=\sum_{y\in{\mathbb{F}}_{2}^{n}}\Big{|}P_{X_{\mathscr{C}}+Z}(y)-\frac{1}{2^{n}}\Big{|}$
$=2d_{\text{\rm TV}}(P_{X_{\mathscr{C}}+Z},P_{U_{n}}).$
The proof is complete. ∎
In the next section, we argue that the reduction procedure can be formulated using the characterization given in this lemma. However, in our calculation, we primarily rely on Definition 2.3 as it offers a deeper insight into the problem.
Remark 1.
Papers [6] and [23] used a slightly different definition of code smoothing, saying that $Z$ smooths a code ${\mathscr{C}}$ if $P_{{\mathsf{G}}Z}\approx P_{U_{k}}$. It is not hard to see that this version of the definition leads to equivalent conclusions.
A characterization of smoothing related to Lemma  2.2 was used to quantify uniformity guarantees for linear hashing in [19]. It was also used in [22] to establish a reduction from LPN to the average-case decoding problem.
The following result is well known.
Lemma 2.3.
Let $e\in{\mathbb{F}}_{2}^{n}$ and $Z\sim P_{Z}$. The bias of the bit $e^{\intercal}Z$ is given by $2^{n-1}\hat{P}_{Z}(e).$
Proof.
We have
$\operatorname{bias}(e^{\intercal}Z)$ $=\frac{\Pr[e^{\intercal}Z=0]-\Pr[e^{\intercal}Z=1]}{2}$
$=\frac{\sum_{z\in{\mathbb{F}}_{2}^{n}}P_{Z}(z){\mathbbm{1}}_{\{e^{\intercal}Z=0\}}-\sum_{z\in{\mathbb{F}}_{2}^{n}}P_{Z}(z){\mathbbm{1}}_{\{e^{\intercal}Z=1\}}}{2}$
$=\frac{\sum_{z\in{\mathbb{F}}_{2}^{n}}P_{Z}(z)(-1)^{e^{\intercal}z}}{2}$
$=2^{n-1}\hat{P}_{Z}(e).\qed$
3. Limitations of the smoothing-based reduction
In this section, we show that the decoding-to-LPN reduction is feasible if the distance to uniformity $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})$ declines slowly as a function of $k$ and is not if it declines faster than $1/\operatorname{poly}(k)$. We start with the impossibility result.
3.1. Positive rate, linear error weight
The main result of this section, Thm. 3.3, provides a bound for $\operatorname{bias}(e^{\intercal}Z)$ in terms of $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})$, error weight $w$, and the dual distance of the code. Using this, we show that for code sequences of growing length $n$ and positive rate, and error weight scaling linearly with $n$, it is impossible to achieve a meaningful reduction if $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,},P_{U_{k}})=\operatorname{negl}(k)$.
Below $B(z,t):=\cup_{i=0}^{t}S(z,i)$ and $V_{n}(t):=|B(z,t)|=\sum_{j=0}^{t}\binom{n}{j}.$ The following two generic lemmas constitute the key technical results that will imply the main theorem.
Lemma 3.1.
Let ${\mathscr{C}}_{0}$ be a code of length $n$ and distance $d$. Let $\varepsilon>0$. Let $\rho$ be a pmf on ${\mathbb{F}}_{2}^{n}$ that satisfies
$d_{\text{\rm TV}}(P_{{\mathscr{C}}_{0}}\ast\rho,P_{U_{n}})\leq\varepsilon.$
Then
$\sum_{x:|x|\leq t}\rho(x)\leq\frac{|{\mathscr{C}}_{0}|V_{n}({t})}{2^{n}}+\varepsilon$
and
$\sum_{x:|x|\geq n-t}\rho(x)\leq\frac{|{\mathscr{C}}_{0}|V_{n}({t})}{2^{n}}+\varepsilon,$
where ${t}=\lfloor\frac{d-1}{2}\rfloor$.
Proof.
Let us prove the first statement. Let ${\mathscr{B}}=\bigcup_{z\in{\mathscr{C}}_{0}}B(z,t)$. Then by (1)
$(P_{{\mathscr{C}}_{0}}\ast\rho)({\mathscr{B}})-P_{U_{n}}({\mathscr{B}})\leq d_{\text{\rm TV}}(P_{{\mathscr{C}}_{0}}\ast\rho,P_{U_{n}})\leq\varepsilon$
Since the balls $B(z,t),z\in{\mathscr{C}}_{0}$ are pairwise disjoint, this implies that
(12) $(P_{{\mathscr{C}}_{0}}\ast\rho)({\mathscr{B}})\leq\frac{V_{n}({t})|{\mathscr{C}}_{0}|}{2^{n}}+\varepsilon.$
Further
$(P_{{\mathscr{C}}_{0}}\ast\rho)({\mathscr{B}})$ $=\sum_{x\in{\mathscr{B}}}\frac{1}{|{\mathscr{C}}_{0}|}\sum_{c\in{\mathscr{C}}_{0}}\rho(x-c)$
$=\frac{1}{|{\mathscr{C}}_{0}|}\sum_{c\in{\mathscr{C}}_{0}}\sum_{x\in{\mathscr{B}}}\rho(x-c)$
$\geq\frac{1}{|{\mathscr{C}}_{0}|}\sum_{c\in{\mathscr{C}}_{0}}\sum_{x\in B(c,t)}\rho(x-c)$
$=\sum_{x:|x|\leq t}\rho(x).$
Combining this with (12) yields the first statement. The proof of the second statement is completely analogous. ∎
Before proceeding to the next lemma, we introduce the following notation, which will be used throughout this section. Let $d^{\bot}$ be a value of the dual distance and let ${t}^{\bot}=\lfloor\frac{d^{\bot}-1}{2}\rfloor$. Let
$I_{n,d^{\bot}}$ =$=\Big{\{}0,1,\dots,{t}^{\bot}\Big{\}}\cup\Big{\{}n-{t}^{\bot},n-{t}^{\bot}+1,\dots,n\Big{\}}$
$I_{n,d^{\bot}}^{\complement}$ =$=\{0,1,\dots,n\}\setminus I_{n,d^{\bot}}.$
Lemma 3.2.
Let $c$ and $C$ be the constants defined in Lemma 2.1. Let ${\mathscr{C}}$ be a code of length $n$ and dual distance $d^{\bot}<n/2$. Let $w$ be an integer satisfying $w\leq cn$. Let $\rho$ be a pmf on ${\mathbb{F}}_{2}^{n}$ satisfying
$d_{\text{\rm TV}}(P_{{\mathscr{C}}^{\bot}}\ast\rho,P_{U_{n}})\leq\varepsilon\quad(\varepsilon>0).$
Then
$\frac{1}{\binom{n}{w}}\sum_{x:|x|=w}2^{n}\hat{\rho}(x)\leq\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n-1}}+Cn\Big{(}1-\frac{2w}{n}\Big{)}^{t^{\bot}}+2\varepsilon.$
Proof.
Using (8) and (9) and viewing ${\mathbbm{1}}_{S(0,w)}$ as a function in the Fourier domain, we have
$\frac{2^{n}}{\binom{n}{w}}\sum_{x:|x|=w}\hat{\rho}(x)$ $=\frac{2^{n}}{\binom{n}{w}}\sum_{x\in{\mathbb{F}}_{2}^{n}}\hat{\rho}(x){\mathbbm{1}}_{S(0,w)}(x)$
$=\frac{1}{\binom{n}{w}}\sum_{y\in{\mathbb{F}}_{2}^{n}}\rho(y)K_{w}(y)$
(13) $\leq\sum_{i\in I_{n,d^{\bot}}}\frac{|K_{w}(i)|}{\binom{n}{w}}\sum_{y:|y|=i}\rho(y)+\sum_{i\in I_{n,d^{\bot}}^{\complement}}\frac{|K_{w}(i)|}{\binom{n}{w}}\sum_{y:|y|=i}\rho(y).$
Let us bound the first sum:
$\sum_{i\in I_{n,d^{\bot}}}\frac{|K_{w}(i)|}{\binom{n}{w}}\sum_{y:|y|=i}\rho(y)$ $\leq\sum_{i\in I_{n,d^{\bot}}}\sum_{y:|y|=i}\rho(y)$
$=\sum_{y:|y|\leq t^{\bot}}\rho(y)+\sum_{y:|y|\geq n-t^{\bot}}\rho(y)$
(14) $\leq\frac{2|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n}}+2\varepsilon,$
where the last inequality is due to Lemma 3.1. Now let us bound the second sum in (13).
From Lemma 2.1 and from the fact that $K_{w}(i)=(-1)^{w}K_{w}(n-i)$, for $w\leq cn$ and $i\in\{0,1,\dots,n\}$ we have
(15) $\frac{|K_{w}(i)|}{\binom{n}{w}}\leq C\Big{(}1-\frac{2w}{n}\Big{)}^{\min\{i,(n-i)\}}.$
Therefore,
$\sum_{i\in I_{n,d^{\bot}}^{\complement}}\frac{|K_{w}(i)|}{\binom{n}{w}}\sum_{y:|y|=i}\rho(y)$ $\leq C\sum_{i\in I_{n,d^{\bot}}^{\complement}}\Big{(}1-\frac{2w}{n}\Big{)}^{\min\{i,(n-i)\}}\sum_{y:|y|=i}\rho(y)$
$\leq C\sum_{i\in I_{n,d^{\bot}}^{\complement}}\Big{(}1-\frac{2w}{n}\Big{)}^{\min\{i,(n-i)\}}$
$\leq C|I_{n,d^{\bot}}|\max_{i\in I_{n,d^{\bot}}^{\complement}}\Big{(}1-\frac{2w}{n}\Big{)}^{\min\{i,(n-i)\}}$
$\leq Cn\Big{(}1-\frac{2w}{n}\Big{)}^{t^{\bot}}.$
Combining this with (14) concludes the proof. ∎
Theorem 3.3.
Let ${\mathscr{C}}$ be an $[n,k,d]$ code with dual distance $d^{\bot}$ and generator matrix ${\mathsf{G}}$. Let $w$ be a non-negative integer satisfying $w\leq cn$. Let $Z\sim P_{Z}$ be a random vector defined on ${\mathbb{F}}_{2}^{n}$ that satisfies $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z},P_{U_{k}})\leq\varepsilon,\varepsilon>0$. Then there exists a vector $e\in{\mathbb{F}}_{2}^{n}$ of weight $w$ such that
(16) $|\operatorname{bias}(e^{\intercal}Z)|\leq\sqrt{\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n+1}}}+\frac{\sqrt{Cn}}{2}\Big{(}1-\frac{2w}{n}\Big{)}^{{t}^{\bot}/2}+\sqrt{\frac{\varepsilon}{2}}.$
Remark 2.
It is not difficult to see that distributions $P_{Z}$ satisfying $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z},P_{U_{k}})\leq\varepsilon$ exist for any $\varepsilon>0$.
Proof.
Setting $\rho=P_{Z}\ast P_{Z}$ and applying the data processing inequality [^2] The data processing inequality applies to any $f$-divergence $D_{f}(P\|Q):={\mathbb{E}}_{Q}\Big{[}f\Big{(}\frac{dP}{dQ}\Big{)}\Big{]}$. In particular, the total variation distance between two measures is an $f$-divergence with $f(s)=\frac{1}{2}|s-1|$. The inequality says that if $P_{X},Q_{X}$ are two measures on $X$, then $D_{f}(P_{X}\|Q_{X})$ decreases if $P_{X}$ and $Q_{X}$ are “passed” through a conditional distribution $P_{Y|X}$ for some set $Y$. In our case $X=Y={\mathbb{F}}_{2}^{n}$, $P_{X}=P_{{\mathscr{C}}^{\bot}}\ast\rho$, $Q_{X}=P_{U_{n}}$, and $P_{Y|X=x}=P_{Z}(\cdot-x)$.
(17) $d_{\text{\rm TV}}(P_{{\mathscr{C}}^{\bot}}\ast\rho,P_{U_{n}})$ $=d_{\text{\rm TV}}(P_{{\mathscr{C}}^{\bot}}\ast P_{Z}\ast P_{Z},P_{U_{n}}\ast P_{Z})\leq d_{\text{\rm TV}}(P_{{\mathscr{C}}^{\bot}}\ast P_{Z},P_{U_{n}}).$
Here we used the fact that $P_{U_{n}}=P_{U_{n}}\ast P_{Z}$ for any distribution $P_{Z}.$ Now, from (11),
$d_{\text{\rm TV}}(P_{{\mathscr{C}}^{\bot}}\ast P_{Z},P_{U_{n}})=d_{\text{\rm TV}}(P_{X_{{\mathscr{C}}^{\bot}+Z}},P_{U_{n}})=d_{\text{\rm TV}}(P_{{\mathsf{G}}Z},P_{U_{k}})\leq\varepsilon.$
Therefore, we have
$d_{\text{\rm TV}}(P_{{\mathscr{C}}^{\bot}}\ast\rho,P_{U_{n}})\leq\varepsilon.$
From Lemma 3.2 we have
(18) $\frac{2^{n}}{\binom{n}{w}}\sum_{x:|x|=w}\hat{\rho}(x)$ $\leq\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n-1}}+Cn\Big{(}1-\frac{2w}{n}\Big{)}^{t^{\bot}}+2\varepsilon.$
From (7), we have $\hat{\rho}(x)=2^{n}\hat{P}_{Z}(x)^{2}$. Therefore,
$\frac{1}{\binom{n}{w}}\sum_{x:|x|=w}(2^{n}\hat{P}_{Z}(x))^{2}$ $\leq\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n-1}}+Cn\Big{(}1-\frac{2w}{n}\Big{)}^{t^{\bot}}+2\varepsilon.$
Set $e=\operatorname*{\arg\!\min}_{x:|x|=w}|2^{n}\hat{P}_{Z}(x)|$. Then
[…]
$\leq\frac{1}{2}\Big{(}\frac{1}{\binom{n}{w}}\sum_{x:|x|=w}|2^{n}\hat{P}_{Z}(x)|\Big{)}$
[…]
$\leq\frac{1}{2}\Big{(}\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n-1}}+Cn\Big{(}1-\frac{2w}{n}\Big{)}^{t^{\bot}}+2\varepsilon\Big{)}^{1/2}$
$=\Big{(}\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n+1}}+\frac{Cn}{4}\Big{(}1-\frac{2w}{n}\Big{)}^{t^{\bot}}+\frac{\varepsilon}{2}\Big{)}^{1/2}$
$\leq\sqrt{\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n+1}}}+\frac{\sqrt{Cn}}{2}\Big{(}1-\frac{2w}{n}\Big{)}^{{t}^{\bot}/2}+\sqrt{\frac{\varepsilon}{2}}.\qed$
Ideally, one would expect to reduce a decoding instance with linear error weight to an LPN problem instance. If this were possible, it would demonstrate that LPN is at least as hard as decoding a generic linear code, which corresponds to an NP-hard problem. However, as we will show next, this is impossible to achieve for linear codes with large dual distance unless $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z},P_{U_{k}})$ is assumed to be very small.
Corollary 3.4.
Let $R\in(0,1)$ and let $\omega<c$, where $c$ is as in Lemma 2.1. For any sequence of $[n_{j},k_{j}]$ linear codes ${\mathscr{C}}_{j},j=1,2,\dots$ of increasing length $n_{j}$ such that ${k_{j}}/{n_{j}}\to R$ and $d({\mathscr{C}}_{j}^{\bot})/n_{j}\to\partial^{\bot}>0$ and any sequence of random vectors $(Z_{j})_{j\geq 1}$ defined on ${\mathbb{F}}_{2}^{n_{j}}$, there exists a sequence of vectors $e_{j}$ with $|e_{j}|/n_{j}\to\omega$ such that the following holds true:
  1. (i)
    If $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{j}Z_{j}},P_{U_{k_{j}}})=\operatorname{negl}(k_{j})$, then $\operatorname{bias}(e_{j}^{\intercal}Z_{j})=\operatorname{negl}(k_{j})$
  2. (ii)
    If $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{j}Z_{j}},P_{U_{k_{j}}})=2^{-\Omega(k_{j})}$, then $\operatorname{bias}(e_{j}^{\intercal}Z_{j})=2^{-\Omega(k_{j})}$,
where ${\mathsf{G}}_{j}$ is the generator matrix of ${\mathscr{C}}_{j}$.
Proof.
The bound on the right-hand side of (16) includes three terms. If the code rate stays positive as $n_{j}$ increases, the Hamming (sphere packing) bound implies that the term $\sqrt{\frac{|{\mathscr{C}}^{\bot}|V_{n}({t}^{\bot})}{2^{n+1}}}$ decays exponentially with $k$. The second term also declines exponentially. Indeed
(19) $\frac{\sqrt{Cn}}{2}\Big{(}1-\frac{2w}{n}\Big{)}^{{t}^{\bot}/2}=O\big{(}\sqrt{n}(1-2\omega)^{\partial^{\bot}n/4}\big{)}=2^{-\Omega(n)},$
where we have omitted the subscripts $j$. Thus the behavior of the bias for $j\to\infty$ is governed by the total variation distance $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{j}Z_{j}},P_{U_{k_{j}}})$ as described in (i), (ii) in the statement. ∎
Remark 3.
(a) Impossibility of reduction in Corollary 3.4 is proved under the assumption of error weight $w$ bounded by $cn$. This is an artifact of the proof rather than an essential limitation. Indeed, the condition $w<cn$ arises because we rely on a simple bound for Krawtchouk polynomials, the estimate (15). Using more refined bounds such as those given in [15], [20, Lemma 2.2] makes it possible to show that a meaningful reduction is unlikely for larger error weights as well. We omit the details.
(b) Just as the restriction on the error weight in Corollary 3.4 is not essential, so is the assumption of the linearly growing dual distance. While typical random linear codes satisfy it with high probability, we can in fact relax it, assuming instead that the dual distance of the codes scales as $d({\mathscr{C}}_{j}^{\bot})=\Omega(n_{j}^{\alpha})$ for some $\alpha\in(0,1)$. Substituting this into (19), we still find that the bias of $e^{\intercal}Z$ declines as $2^{-\Omega(n^{\alpha})}$. We note that there are abundantly many code sequences with dual distance satisfying these constraints, for instance, with high probability random codes are among them.
We close this section with a remark concerning the range of parameters for the LPN reduction. We have shown that the assumption $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{j}Z_{j}},P_{U_{k_{j}}})=\operatorname{negl}(k)$ makes meaningful reductions with linear error weight impossible. This leaves the option of reducing decoding to LPN with errors of sublinear weight. We argue that if the weight satisfies $w=o(k)$, where $k$ is the dimension of the code, then a meaningful reduction is possible. At the same time, as noted earlier, this implication does not contribute toward the desired hardness proof of LPN.
According to Theorem 3.3, given that $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z},P_{U_{k}})=\operatorname{negl}(k)$, a necessary condition for $\operatorname{bias}(e^{\intercal}Z)=\Omega(1/\operatorname{poly}(k))$ is
$\frac{\sqrt{Cn}}{2}\Big{(}1-\frac{2w}{n}\Big{)}^{{t}^{\bot}/2}=\Omega(1/\operatorname{poly}(k)).$
For codes whose dual distance $d^{\bot}$ scales linearly with the block length, the above condition is equivalent to $w=O(\log(k))$. We state this as another corollary whose proof is immediate.
Corollary 3.5.
Let $R,\partial^{\bot}\in(0,1)$, and $l\in{\mathbb{Z}}^{+}$. Let ${\mathscr{C}}_{j}$ be a sequence of $[n_{j},k_{j}]$ linear codes of increasing length $n_{j}$ such that ${k_{j}}/{n_{j}}\to R$ and $d({\mathscr{C}}_{j}^{\bot})/{n_{j}}\to\partial^{\bot}$. Denote by ${\mathsf{G}}_{j}$ the generator matrix of ${\mathscr{C}}_{j}$. For all $j$, let $e_{j}\in{\mathbb{F}}_{2}^{n_{j}}$ be a vector satisfying $|e_{j}|=w_{j}$. If $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{j}Z_{j}},P_{U_{k_{j}}})=\operatorname{negl}(k_{j})$, then the following holds true:
[…]
It is interesting to note that the result of Corollary 3.5 is essentially tight. Namely, assuming Bernoulli noise, Proposition 5.1 in [8] implies that $w=\Theta(\log(k))$ is sufficient to achieve a meaningful reduction.
3.2. Slow Smoothing allows Reduction
In Section 3.1, we studied the case where the error weight scales linearly with the block length, with the additional assumption that $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{j}Z_{j}},P_{U_{k_{j}}})$ decays fast. Here we show that lifting this assumption makes it possible to obtain a meaningful reduction.
Theorem 3.6.
Let $R\in(0,1)$, $\omega\in(0,1/2)$, and $l\in{\mathbb{Z}}^{+}$. Let ${\mathscr{C}}_{j}$ be a sequence of $[n_{j},k_{j}]$ linear codes of increasing length $n_{j}$ such that ${k_{j}}/{n_{j}}\to R$. Let ${\mathsf{G}}_{j}$ be a generator matrix of ${\mathscr{C}}_{j}$ and let $e_{j}\in{\mathbb{F}}_{2}^{n_{j}}$ be a vector satisfying $|e_{j}|=\lfloor\omega n_{j}\rfloor$. Then there exists a sequence of distributions $P_{Z_{j}}$ satisfying the following conditions:
  1. (i)
    $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{j}Z_{j},e_{j}^{\intercal}Z_{j}},P_{U_{k_{j}}}P_{e_{j}^{\intercal}Z_{j}})=O(k_{j}^{-l})$,
  2. (ii)
    $\operatorname{bias}(e_{j}^{\intercal}Z_{j})=\Omega(k_{j}^{-l})$.
This theorem follows directly from the next lemma upon setting $\gamma=k^{-l}$.
Lemma 3.7.
Let ${\mathscr{C}}\in{\mathbb{F}}_{2}^{n}$ be linear code satisfying $|{\mathscr{C}}|<{2^{n-1}}/V_{n}(2)$ and let $e\in{\mathbb{F}}_{2}^{n}\setminus\{0\},|e|<n/2$. Then for any $\gamma>0$, the following conditions are simultaneously achievable:
  1. (i)
    $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,e^{\intercal}Z},P_{U_{k}}P_{e^{\intercal}Z})\leq\gamma\big{(}\frac{5}{2}-\frac{|e|}{n}\big{)}$,
  2. (ii)
    $\operatorname{bias}(e^{\intercal}Z)=\frac{\gamma}{2}\big{(}1-\frac{2|e|}{n}\big{)}$.
Proof.
Let $P_{Z}=(1-\gamma)P_{U_{n}}+\gamma\frac{{\mathbbm{1}}_{S(0,1)}}{n}$. By Lemma 2.3 and the definition of $K_{w}(i)$
$\operatorname{bias}(e^{\intercal}Z)=2^{n-1}\hat{P}_{Z}(e)=\frac{1}{2}\Big{(}(1-\gamma){\mathbbm{1}}_{\{0\}}(e)+\frac{\gamma}{n}K_{1}(|e|)\Big{)}=\frac{\gamma}{2}\Big{(}1-\frac{2|e|}{n}\Big{)}.$
[…]
$d_{\text{\rm TV}}(P_{{\mathsf{G}}_{e}Z},P_{U_{k}}P_{e^{\intercal}Z})\leq d_{\text{\rm TV}}(P_{U_{k+1}},P_{U_{k}}P_{e^{\intercal}Z})+d_{\text{\rm TV}}(P_{{\mathsf{G}}_{e}Z},P_{U_{k+1}}).$
To bound the first term, use (2) to write
$d_{\text{\rm TV}}(P_{U_{k+1}},P_{U_{k}}P_{e^{\intercal}Z})$ $=\frac{1}{2}\sum_{x\in{\mathbb{F}}_{2}^{k},b\in{\mathbb{F}}_{2}}|P_{U_{k+1}}(x,b)-P_{U_{k}}(x)P_{e^{\intercal}Z}(b)|$
$=\frac{1}{2}\sum_{x\in{\mathbb{F}}_{2}^{k}}\big{|}\frac{1}{2^{k+1}}-\frac{1}{2^{k}}P_{e^{\intercal}Z}(0)\big{|}+\frac{1}{2}\sum_{x\in{\mathbb{F}}_{2}^{k}}\big{|}\frac{1}{2^{k+1}}-\frac{1}{2^{k}}P_{e^{\intercal}Z}(1)\big{|}$
$=\operatorname{bias}(e^{\intercal}Z)$
$=\frac{\gamma}{2}\big{(}1-\frac{2|e|}{n}\big{)}.$
Now we are left to bound $d_{\text{\rm TV}}(P_{{\mathsf{G}}_{e}Z},P_{U_{k+1}})$. From (11), we have
$d_{\text{\rm TV}}(P_{{\mathsf{G}}_{e}Z},P_{U_{k+1}})=d_{\text{\rm TV}}(P_{X_{{\mathscr{C}}_{e}}+Z},P_{U_{n}}).$
We divide the analysis into two cases depending on the value of $d^{\bot}({\mathscr{C}}_{e})$. First assume that $d({\mathscr{C}}_{e}^{\bot})\geq 3$. In this case, we obtain
$P_{X_{{\mathscr{C}}_{e}^{\bot}}+Z}(x)=(1-\gamma)P_{U_{n}}(x)+\gamma\Big{(}\frac{{\mathbbm{1}}_{S(0,1)}}{n}\ast P_{{\mathscr{C}}_{e}^{\bot}}\Big{)}(x)=\frac{1-\gamma}{2^{n}}+\frac{\gamma{\mathbbm{1}}_{{\mathcal{S}}}(x)}{n|{\mathscr{C}}_{e}^{\bot}|},$
where ${\mathcal{S}}=\bigcup_{z\in{\mathscr{C}}_{e}^{\bot}}S(z,1)$. Therefore,
$d_{\text{\rm TV}}(P_{X_{{\mathscr{C}}_{e}^{\bot}}+Z},P_{U_{n}})$ $=\frac{1}{2}\sum_{x\in{\mathbb{F}}_{2}^{n}}\Big{|}(P_{Z}\ast P_{{\mathscr{C}}_{e}^{\bot}})(x)-\frac{1}{2^{n}}\Big{|}$
$=\sum_{x\in{\mathcal{S}}}\Big{|}\frac{1-\gamma}{2^{n}}+\frac{\gamma}{n|{\mathscr{C}}_{e}^{\bot}|}-\frac{1}{2^{n}}\Big{|}+\sum_{x\in{\mathbb{F}}_{2}^{n}\setminus{\mathcal{S}}}\Big{|}\frac{1-\gamma}{2^{n}}-\frac{1}{2^{n}}\Big{|}$
$=n|{\mathscr{C}}_{e}^{\bot}|\gamma\Big{(}\frac{1}{n|{\mathscr{C}}_{e}^{\bot}|}-\frac{1}{2^{n}}\Big{)}+(2^{n}-n|{\mathscr{C}}_{e}^{\bot}|)\frac{\gamma}{2^{n}}$
$=2\gamma\Big{(}1-\frac{n|{\mathscr{C}}_{e}^{\bot}|}{2^{n}}\Big{)}$
$\leq 2\gamma.$
Now assume that $d({\mathscr{C}}_{e}^{\bot})\in\{1,2\}$. Since $|{\mathscr{C}}_{e}|\leq 2^{n}/V_{n}(2)$, $|{\mathscr{C}}_{e}^{\bot}|>V_{n}(2)$, so there exists a nontrivial subcode ${\mathscr{C}}^{\prime}\subset{\mathscr{C}}_{e}^{\bot}$ such that $d({\mathscr{C}}^{\prime})\geq 3$. By the above argument we have
$d_{\text{\rm TV}}(P_{X_{{\mathscr{C}}^{\prime}}+Z},P_{U_{n}})\leq 2\gamma.$
Let $\mathcal{L}$ be the set of coset leaders in ${\mathscr{C}}_{e}^{\bot}/{\mathscr{C}}^{\prime}$. Then we can write $P_{{\mathscr{C}}_{e}^{\bot}}=\frac{1}{|{\mathscr{C}}_{e}^{\bot}/{\mathscr{C}}^{\prime}|}\sum_{l\in\mathcal{L}}P_{{\mathscr{C}}^{\prime}+l}$. From the convexity of the total variation distance, we have
$d_{\text{\rm TV}}(P_{X_{{\mathscr{C}}_{e}^{\bot}}+Z},P_{U_{n}})\leq\frac{1}{|\mathcal{L}|}\sum_{l\in\mathcal{L}}d_{\text{\rm TV}}(P_{X_{{\mathscr{C}}^{\prime}+l}+Z},P_{U_{n}})\leq 2\gamma.$
Therefore, $d_{\text{\rm TV}}(P_{{\mathsf{G}}Z,e^{\intercal}Z},P_{U_{k}}P_{b})\leq 2\gamma+\frac{\gamma}{2}\big{(}1-\frac{2|e|}{n}\big{)}=\gamma\big{(}\frac{5}{2}-\frac{|e|}{n}\big{)}$. Since $|e|<n/2$, in both cases condition (i) is satisfied. Thus, the proof is complete. ∎
Acknowledgments
Research supported by NSF grants CCF-2110113 (NSF-BSF), CCF-2104489, and CCF-2330909. This work was partly done when the authors were visiting the Simons Institute for the Theory of Computing in Spring 2024. The authors are grateful to Nicolas Resch for a useful discussion of the decoding-to-LPN reduction problem.